# Information Theory for Complex Systems Scientists: What, Why, & How?

### v3

@@ Abstract
In the 21st century, many of the crucial scientific and technical issues facing humanity can be understood as problems associated with understanding, modeling, and ultimately controlling complex systems: systems comprised of a large number of non-trivially interacting components whose collective behavior can be difficult to predict. Information theory, a branch of mathematics historically associated with questions about encoding and decoding messages, has emerged as something of a lingua franca for those studying complex systems, far exceeding its original narrow domain of communication systems engineering. In the context of complexity science, information theory provides a set of tools which allow researchers to uncover the statistical and effective dependencies between interacting components; relationships between systems and their environment; mereological whole-part relationships; and is sensitive to non-linearities missed by commonly parametric statistical models.
In this review, we aim to provide an accessible introduction to the core of modern information theory, aimed specifically at aspiring (and established) complex systems scientists. This includes standard measures, such as Shannon entropy, relative entropy, and mutual information, before building to more advanced topics, including: information dynamics, measures of statistical complexity, information decomposition, and effective network inference. In addition to detailing the formal definitions, in this review we make an effort to discuss how information theory can be interpreted and develop the intuition behind abstract concepts like ”entropy,” in the hope that this will enable interested readings to understand what information is, and how it is used, at a more fundamental level.
August 13, 2024
@@ I What Are Complex Systems?
If the first decades of ${}^{\text{th}}$ century have provided any insight into what the scientific challenges of tomorrow will be, it is that many of the core problems are tied together by a common thread: complexity emerging from highly-interconnected systems. From climate change, to pandemics, from disinformation to economic inequality, a defining characteristic of many issues facing the modern world is that they deal with the problem of modeling, predicting, and controlling emergent properties of systems comprised of thousands, or millions of interacting elements. In the case of pandemics, those elements are individual humans and animals, milling and seething across the globe. They spread germs, get sick, get better, and occasionally die. In the case of “fake news”, the elements form an ecosystem of humans and ever-more human-like artificial intelligences, all navigating online environments controlled by increasingly opaque algorithms that were trained on tremendous amounts of data. Climate change threatens the stability of a huge number of interconnected, complex systems, from coral reef ecosystems (which bleach in the face of sustained high temperatures [1]), to agriculture and food-production systems (possibly leading to food shocks, and in extreme cases, famine [2]).
These problems of complexity are interconnected as well: far from being restricted to the traditional disciplinary silos of Academia (physics, chemistry, biology, etc), complex systems can involve many different areas of science simultaneously and require seeing connections across domains. For example, it takes tremendous amounts of energy to train the artificial intelligences that increasingly dominate our online lives. In an economy run on fossil fuel, this means that the systems driving one phenomenon (“fake news”, social media bubbles, “infodemics” [3]) are simultaneously exacerbating another one (climate change) with their large carbon footprints and outrageous water demands [4, 5, 6]. These simple examples highlight some of the fundamental challenges of modern complex systems science:
@@
    All of the systems described above are highly non-linear [7], exhibiting complex phenomena such as critical phase changes [8], cascading failures [9], extreme ranges of intensities [10], and self-organization [11] (to name but a few of the interesting properties complex systems can display).
@@
    These systems are multiscale: at the fundamental level, they involve millions, or even billions of moving parts: modeling the entire structure would require astronomical amounts of data, computing power, and very detailed mathematical and/or algorithmic models of behavior. Consequently, the ability to think critically about the distinction between “wholes” and “parts” is key [12]. So is understanding when it is possible to reduce the dimensionality with things like coarse-graining [13]. Furthermore, “higher-order” or “synergistic” effects present only at the macro-scale can complicate analyses when systems don’t neatly decompose into sums of parts [14].
@@
    The processes that make up the world are dynamic: the past informs on the future and individual moments cannot be assumed to be decoupled from the history that lead up to that point [15].
What might be called “classical” statistical modelling, which makes assumptions of linearity, reductionism, and independence can struggle when attempting to disentangle complex systems. Warren Weaver, a pioneer of early complex system science described a schema with which systems can be categorized:
@@
What, then, is the way forward? Science in the 21st century needs a toolkit that can be applied to as many complex systems as possible, both to understand their inner workings, and facilitate interdisciplinary interactions between fields that may never have had much reason to come into contact with each-other, and therefore lack a common language. Furthermore, rather than running from the nonlinear, hard-to-model nature of complex data, the science of tomorrow must be able to tackle these systems without throwing out important information by making over-simplifying assumptions of linearity, independence, or equilibrium conditions. Increasingly, the emerging answer to this is “information theory,” which is poised to become the lingua franca of our increasingly complexity-aware scientific enterprise.
@@ II What Is Information?
Information theory was first developed by Claude Shannon, an engineer working at Bell Labs on the problem of how to reliably send a signal from a sender to a receiver over some potentially noisy channel (such as a telegraph line). In a now-famous moment of remarkable genius, Shannon produced almost the entirety of classical information theory in a single monograph: A Mathematical Theory of Communication, first published in 1948 [17]. In this highly practical context, “information” was understood as referring to signals being propagated through a channel. We will describe the idea of information in formal detail below, but intuitively, it can be understood in terms of how our uncertainly is resolved by the data made available to us. If someone speaks into a telephone receiver, but all the person at the other end hears is static white-noise, the “information” has clearly been lost: the listener has no way of inferring what the speaker said. In contrast, if the voice comes through loud and clear, then the listener has no uncertainty about what the speaker said, and information has been transmitted. Shannon provided a mathematically rigorous formal language with which to discuss how one might quantify how much information was transmitted (or lost), and what the fundamental limits on transmission are in noisy environments. [^1] It is not hyperbolic to say that this work formed one of the fundamental foundations on which our modern, digital world was built (on par with the development of electricity, or the internal combustion engine in terms of impact).
@@
The significance of this development was not lost on Shannon’s fellow cyberneticians, and information theory quickly generated interest from a broad and interdisciplinary group of scientists, mathematicians, and even philosophers. The question of what information theory’s “natural domain” is has been a contentious one ever since. Despite Shannon’s initial narrow, and eminently practical focus, during the cybernetic revolution in ${}^{\text{th}}$ information theory was quickly applied to a wide range of fields, from sociology to biology [18, 19]. Shannon himself, alarmed at what he saw as a the hasty misapplication of his theories to disciplines where they didn’t belong, pushed back, claiming that, by the mid-1950s that information theory had become a “bandwagon” and was been “oversold” by overeager researchers [20]. Despite Shannon’s apparent misgivings, however, in the modern day, information theory has retained its popularity, finding applications in diverse fields such as neuroscience [21, 22], ecology [23, 24], artificial intelligence [25, 26], sociology [27], climatology [28, 29, 30, 31], and deep connections have been found between information theoretic formalisms and fundamental physical models such as thermodynamics [32, 33]. Finally, information theory has become a key foundation on which a variety of formal theories of “emergence” have been constructed [34, 35, 36, 37, 38, 39, 40, 41]. Emergence, the phenomenon where a system collectively displays properties not trivially reducible to some property of its component parts is arguably the defining feature of complex systems, and the links between multivariate information theory and emergence solidify the deep link between the fields.
@@
It is interesting to consider the question of how it came to be that Shannon’s theory, developed within the practical confines of communications engineering, has been able to find such widespread success far beyond its original home. Has the bandwagon simply continued to run amok, heedless of Shannon’s 1956 warning? I would (perhaps unsurprisingly) argue ”no” (it’s hard to imagine why I would have written this manifesto otherwise). Rather, I claim that information theory’s native domain is not just the field of communication engineering where it was born, but rather information theory is better understood as being fundamentally about the general mathematics of inference [26], particularly, inference under conditions of uncertainty. The original applications of information theory (how to reliably transmit a message from a sender to a receiver along a noisy channel) is itself a particular kind of inference problem; a special case of a more general framework. When the receiver receives a message transmitted in the presence of noise, the problem of decoding that message is one of making inferences about the most probable contents of the original message (“what was the sender likely saying”). Similarly, when engineering a usable communication channel, the goal is to give the receiver the best possible chance of correctly inferring the senders message. It is the logic of inference (encoding, decoding, probability and statistics) that informs the development of communications channels, not the other way around.
@@
This perspective of “information theory as inference” resolves another long-standing tension in the field. Information theory is agnostic to the question of the “meaning” of a message: it never mattered to Shannon what message was being sent over a noisy channel, only the statistics of the symbols that make it up. This distinction has been referred to as the difference between the syntactic information (which information theory is concerned with), and the semantic information (which it is not) [42]. As long as information theory has been understood as a “mathematical theory of communication”, the question of meaning (semantic content) is a very natural one: everyone cares more about the meaning of a message than the statistics of the symbols used to transmit it  [^2] Except engineers, of course and information theory’s failure to address this has lead some to doubt its appropriateness as a theory of communication.
@@
This paper is written with perspective that focusing on the “communication” angle is a mistake, or at least, insufficiently general. When considering “information theory as inference”, it is clear that the question of semantic information is largely irrelevant: we are not interested in what a message might mean, but rather, ensuring that we have inferred the correct message from some space of possible messages. The same mathematics, however can apply to any issue where we are attempting to increase our degree of certainty about the state of some as-yet-unknown variable. Whether it is a message coming down a pipe, the future state of an evolving system, or the degree of mutual coupling between two variables. These are the core questions that go into modelling complex systems.
@@ A Note to the Mathematically Nervous
It is impossible to write a comprehensive review of information theory (a field of mathematics in it’s own right) in complex systems (a very quantitative meta-field) without relying on a certain amount of mathematics. However, throughout this review we have done our best to bolster the formulae with discussions about the intuition behind the measures and how they might be interpreted. Complex systems draws interest from many fields, and many (especially students just discovering the interest for the first time) may not have a comprehensive mathematics background. To paraphrase Albert Einstein, we have tried to “make everything as simple as possible, but no simpler”. For readers who wish the refresh on the basics on probability theory (on which information theory is fundamentally based), see the Supplementary Information).
@@ II.1 Entropy
    How uncertain am I about the state of some variable?
@@
    How much “work” (for want of a better phrase) do I need to do to become more certain than I am now?
It is important to note that these two core questions are not (on the surface) actually about information, but rather, they are about uncertainty, and how uncertainty changes. This forms the core of Shannon’s great insight: information is the reduction of uncertainty associated with observing data. To even begin to mathematically understand information, we must first understand uncertainty. This is done with the “entropy”, a mathematical function that quantifies how uncertain we are about the state of some variable.
There is a (likely apocryphal) story that, upon developing his theory, Claude Shannon approached the famous physicist and fellow cybernetician John von Neumann and asked what he should call his new measure. Von Neumann (apparently) replied: “you should call it entropy…since no one really knows what entropy is, you’ll always have the upper hand in a debate.” (Paraphrased). While tongue-in-cheek, this story highlights the reality that, despite Shannon’s originally narrow focus, the structure he developed is remarkably general and lends itself to many interpretations (and more than a few misinterpretations as well).
@@ II.1.1 Entropy as Expected Surprise
To answer Question 1 (”how uncertain am I about the state of something in the world”), we need a measure of uncertainty; a way to quantify exactly how certain we are about whatever it is we believe. To understand what we mean by quantifying “uncertainty”, consider the case of a fair die, where the probability of any face coming up is 1/6. If you were gambling on the die, you would have no reason to pick any face over another and consequently no outcome would be more surprising than any other. You would be no more shocked to roll a 2 than a 5 and are, in some sense, maximally uncertain about the outcome of a roll. This notion of “surprise” as being related to “uncertainty” becomes clearer if we instead imagine that you are a cheater and, at some point, manage to switch the fair die for a loaded one of your own. This tricky die is balanced in such a way that 2 comes up 2/3 of the time, and all other faces come up with probability 1/15. Knowing this probability distribution ahead of time, you will be less surprised to see a 2 come up than you would be to see a 2 come up on the fair die. Similarly, you would be more surprised to see any other number turn up on the loaded die than you would to see that same number on the fair die. In this case, you are more certain about the outcome of a roll than you would be in the case of the fair die.
Everyone has an intuitive notion that some events are more surprising than others, and that the amount of surprise you experience is directly related to how likely you were to make a correct prediction. You’d be more surprised to receive a spaceship for Christmas than a pair of socks, because socks are a much more likely gift than spaceships! This formed the basis of Shannon’s derivation entropy, and is the first, fundamental mathematical measure we will introduce. Shannon [17] realized that a quantitative measure of surprise (noted as $h()$) should satisfy a set of four basic criteria. For a random variable $X$ that draws states from the support set $\mathcal{X}$ according to the probability distribution $P(x)$:

### v4

@@ Abstract
In the 21st century, many of the crucial scientific and technical issues facing humanity can be understood as problems associated with understanding, modeling, and ultimately controlling complex systems: systems comprised of a large number of non-trivially interacting components whose collective behavior can be difficult to predict. Information theory, a branch of mathematics historically associated with questions about encoding and decoding messages, has emerged as something of a lingua franca for those studying complex systems, far exceeding its original narrow domain of communication systems engineering. In the context of complexity science, information theory provides a set of tools which allow researchers to describe a variety of dependencies, including interactions between the component parts of a system, interactions between a system and it’s environment, and the mereological interaction between the parts and the “whole”.
In this review, we aim to provide an accessible introduction to the core of modern information theory, aimed specifically at aspiring (and established) complex systems scientists. This includes standard measures, such as Shannon entropy, relative entropy, and mutual information, before building to more advanced topics, including: information dynamics, measures of statistical complexity, information decomposition, and effective network inference. In addition to detailing the formal definitions, we also make an effort to discuss how information theory can be interpreted and to develop the intuitions behind abstract concepts like “entropy”. We aim to enable interested readers to understand what information is, and how it is used, at a fundamental level to better further their own research and education.
January 17, 2025
@@ I What Are Complex Systems?
If the first decades of ${}^{\text{th}}$ century have provided any insight into what the scientific challenges of tomorrow will be, it is that many of the core problems are tied together by a common thread: complexity emerging from highly-interconnected systems. From climate change, to pandemics, from disinformation to economic inequality, a defining characteristic of many issues facing the modern world is that they deal with the problem of modeling, predicting, and controlling emergent properties of systems comprised of thousands, or millions of interacting elements. In the case of pandemics, those elements are individual humans and animals, milling and seething across the globe. Animals and humans interact, and in doing so, spread evolving pathogens; they fall ill, thy recover, and occasionally, they die. In the context of “fake news”, the elements form an ecosystem of humans and ever-more human-like artificial intelligences, all navigating online environments controlled by increasingly opaque algorithms. Climate change threatens the stability of a huge number of interconnected, complex systems, from coral reef ecosystems (which bleach in the face of sustained high temperatures [1]), to agriculture and food-production systems (possibly leading to food shocks, and in extreme cases, famine [2]).
These problems of complexity are interconnected as well: far from being restricted to the traditional disciplinary silos of Academia (physics, chemistry, biology, etc), complex systems can involve many different areas of science simultaneously, and require seeing connections across domains. For example, it takes tremendous amounts of energy to train the artificial intelligences that increasingly dominate our online lives. In an economy run on fossil fuel, this means that the systems driving one phenomenon (“fake news”, social media bubbles, “infodemics” [3]) are simultaneously exacerbating another one (climate change) with their large carbon footprints and outrageous water demands [4, 5, 6]. These simple examples highlight some of the fundamental challenges of modern complex systems science:
@@
    All of the systems described above are highly non-linear [7], exhibiting complex phenomena such as critical phase changes (or “tipping points”) [8], cascading failures [9], extreme ranges of intensities [10], and self-organization [11] (to name but a few of the interesting properties complex systems can display).
@@
    These systems are multiscale: at the fundamental level, they involve millions, or even billions of moving parts: modeling the entire structure would require astronomical amounts of data, computing power, and very detailed mathematical and/or algorithmic models of behavior. Consequently, the ability to think critically about the distinction between “wholes” and “parts” is key [12]. So is understanding when it is possible to reduce the dimensionality via processes such as coarse-graining [13]. Furthermore, “higher-order” or “synergistic” effects present only at the macro-scale can complicate analyses when systems don’t neatly decompose into sums of parts [14].
@@
    The processes that make up the world are dynamic: the past informs on the future and individual moments cannot be assumed to be decoupled from the history that lead up to that point [15]. This can take the form of chaotic sensitive dependence on initial conditions, dynamic hysteresis, or long-term autocorrelations (also called “memory”), all of which can describe the influence of the past on the future.
What might be called “classical” statistical modeling, which makes assumptions of linearity, reductionism, and independence can struggle when attempting to disentangle complex systems. Warren Weaver, a pioneer of early complex system science described a schema with which systems can be categorized:
@@
What, then, is the way forward? Science in the 21st century needs a toolkit that can be applied to as many complex systems as possible, both to understand their inner workings, and facilitate interdisciplinary interactions between fields that may never have had much reason to come into contact with each-other, and therefore lack a common language. Furthermore, rather than running from the nonlinear, hard-to-model nature of complex data, the science of tomorrow must be able to tackle these systems without throwing out important information or making over-simplifying assumptions of linearity, independence, or equilibrium conditions. Increasingly, the emerging answer to this problem is “information theory,” which is poised to become the lingua franca of our increasingly complexity-aware scientific enterprise.
@@ II What Is Information?
Information theory was first developed by Claude Shannon, an engineer working at Bell Labs on the problem of how to reliably transmit a signal from a sender to a receiver over some potentially noisy channel (such as a telegraph line). In a now-famous moment of remarkable genius, Shannon produced almost the entirety of classical information theory in a single monograph: A Mathematical Theory of Communication, first published in 1948 [17]. In this highly practical context, “information” was understood as referring to signals being propagated through a channel. We will describe the idea of information in formal detail below, but intuitively, it can be understood in terms of how our uncertainly is resolved by the data made available to us. If someone speaks into a telephone receiver, but all the person at the other end hears is static white-noise, the “information” has clearly been lost: the listener has no way of inferring what the speaker said. In contrast, if the voice comes through loud and clear, then the listener has no uncertainty about what the speaker said, and information has been transmitted. If the channel is noisy, but not totally compromised (perhaps the sound quality is poor), the listener may not be able to perfectly catch every word, but can still reconstruct the gist of the message based on the parts that they could hear. Shannon provided a mathematically rigorous formal language with which one might quantify how much information was transmitted (or lost), and what the fundamental limits on transmission are in noisy environments. [^1] It is not hyperbolic to say that this work formed one of the fundamental foundations on which our modern, digital world was built (on par with the development of electricity, or the internal combustion engine in terms of impact).
@@
The significance of this development was not lost on Shannon’s fellow cyberneticians, and information theory quickly received interest from a broad and interdisciplinary group of scientists, mathematicians, and even philosophers. The question of what information theory’s “natural domain” is has been a contentious one ever since. Despite Shannon’s initial narrow, and eminently practical focus, during the cybernetic revolution of ${}^{\text{th}}$ information theory was quickly applied to a wide range of fields, from sociology to biology [18, 19]. Shannon himself, alarmed at what he saw as a the hasty misapplication of his theories to disciplines where they didn’t belong, pushed back, claiming that, by the mid-1950s that information theory had become a “bandwagon” and was been “oversold” by overeager researchers [20]. Despite Shannon’s apparent misgivings, however, in the modern day, information theory has retained its popularity, finding applications in diverse fields such as neuroscience [21, 22], ecology [23, 24], artificial intelligence [25, 26], sociology [27], climatology [28, 29, 30, 31], and deep connections have been found between information theoretic formalisms and fundamental physical models such as thermodynamics (forming the basis of “stochastic thermodynamics” [32]). Finally, information theory has become a key foundation on which a variety of formal theories of “emergence” have been constructed [33, 34, 35, 36, 37, 38, 39, 40]. Emergence, the phenomenon where a system collectively displays properties not trivially reducible to some property of its component parts is arguably the defining feature of complex systems, and the links between multivariate information theory and emergence solidify the deep link between the fields.
@@
It is interesting to consider the question of how it came to be that Shannon’s theory, developed within the practical confines of communications engineering, has been able to find such widespread success far beyond its original home. Has the bandwagon simply continued to run amok, heedless of Shannon’s 1956 warning? I would (perhaps unsurprisingly) argue “no” (it’s hard to imagine why I would have written this tutorial review otherwise). Rather, I claim that information theory’s native domain is not just the field of communication engineering where it was born, but rather information theory is better understood as being fundamentally about the general mathematics of inference [26], particularly, inference under conditions of uncertainty. The original applications of information theory (how to reliably transmit a message from a sender to a receiver along a noisy channel) is itself a particular kind of inference problem; a special case of a more general framework. When the receiver receives a message transmitted in the presence of noise, the problem of decoding that message is one of making inferences about the most probable contents of the original message (“what was the sender likely saying”). Similarly, when engineering a usable communication channel, the goal is to give the receiver the best possible chance of correctly inferring the senders message. It is the logic of inference (encoding, decoding, probability and statistics) that informs the development of communications channels, not the other way around.
@@
This perspective of “information theory as inference” resolves another long-standing tension in the field. Information theory is agnostic to the question of the “meaning” of a message: it never mattered to Shannon what message was being sent over a noisy channel, only the statistics of the symbols that make it up. This distinction has been referred to as the difference between the syntactic information (which information theory is concerned with), and the semantic information (which it is not) [41]. As long as information theory has been understood as a “mathematical theory of communication”, the question of meaning (semantic content) is a very natural one: everyone cares more about the meaning of a message than the statistics of the symbols used to transmit it  [^2] Except engineers, of course. and information theory’s failure to address this has lead some to doubt its appropriateness as a theory of communication.
@@
This paper is written with the perspective that focusing on the “communication” angle is a mistake, or at least, insufficiently general. When considering “information theory as inference”, it is clear that the question of semantic information is largely irrelevant: we are not interested in what a message might mean, but rather, ensuring that we have inferred the correct message from some space of possible messages. The same mathematics, however can apply to any issue where we are attempting to increase our degree of certainty about the state of some as-yet-unknown variable. Whether it is a message coming down a pipe, the future state of an evolving system, or the degree of mutual coupling between two variables. These are the core questions that form the basic problem of modeling complex systems in any domain.
@@ A Note to the Mathematically Nervous
It is impossible to write a comprehensive review of information theory (a field of mathematics in its own right) in complex systems (a very quantitative meta-field) without relying on a certain amount of mathematics. However, throughout this review I have done my best to bolster the formulae with discussions about the intuitions behind the math, and how the measures might be interpreted. Complex systems draws interest from many fields, and many (especially students just discovering the interest for the first time) may not have a comprehensive mathematics background. To paraphrase Albert Einstein, we have tried to “make everything as simple as possible, but no simpler”. For readers who wish to refresh on the basics of probability theory (on which information theory is fundamentally based), see the Supplementary Information).
@@ II.1 Entropy
    How uncertain am I about the state of some thing I could observe?
@@
    How much “work” (for want of a better word) do I need to do to become more certain than I am now?
It is important to note that these two core questions are not (on the surface) actually about information, but rather, they are about uncertainty, and how uncertainty changes. This forms the core of Shannon’s great insight: information is the reduction of uncertainty associated with observing data. To even begin to mathematically understand information, we must first understand uncertainty. This is done with the “entropy”, a mathematical function that quantifies how uncertain we are about the state of some variable given its statistics.
There is a (likely apocryphal) story that, upon developing his theory, Claude Shannon approached the famous physicist and fellow cybernetician John von Neumann and asked what he should call his new measure. Von Neumann (allegedly) replied: “you should call it entropy…since no one really knows what entropy is, you’ll always have the upper hand in a debate.” (Paraphrased). While tongue-in-cheek, this story highlights the reality that, despite Shannon’s originally narrow focus, the structure he developed is remarkably general and lends itself to many interpretations (and more than a few misinterpretations as well).
@@ II.1.1 Entropy as Expected Surprise
To answer Question 1 (”how uncertain am I about the state of something in the world”), we need a measure of uncertainty; a way to quantify exactly how certain we are about whatever it is we believe. To understand what we mean by quantifying “uncertainty”, consider the case of a fair die, where the probability of any face coming up is 1/6. If you were gambling on the die, you would have no reason to pick any face over another and consequently no outcome would be more surprising than any other. You would be no more shocked to roll a 2 than a 5 and are, in some sense, maximally uncertain about the outcome of a roll. This notion of “surprise” as being related to “uncertainty” becomes clearer if we instead imagine that you are a cheater and, at some point, managed to switch the fair die for a loaded one of your own. This tricky die is balanced in such a way that 2 comes up 2/3 of the time, and all other faces come up with probability 1/15. Knowing this probability distribution ahead of time, you will be less surprised to see a 2 come up than you would be to see a 2 come up on the fair die. Similarly, you would be more surprised to see any other number turn up on the loaded die than you would to see that same number on the fair die. In this case, you are more certain about the outcome of a roll with the tricky die than you would be in the case of the fair one.
Everyone has an intuitive notion that some events are more surprising than others, and that the amount of surprise you experience is directly related to how likely you were to make a correct prediction. You’d be more surprised to receive a spaceship for Christmas than a pair of socks, because socks are a much more likely gift than spaceships! This formed the basis of Shannon’s derivation of the entropy, and is the first, fundamental mathematical measure we will introduce. Shannon [17] realized that a quantitative measure of surprise (noted as $h()$) should satisfy a set of four basic criteria. For a random variable $X$ that draws states from the support set $\mathcal{X}$ according to the probability distribution $P(x)$: